### v1

# Carolina Torreblanca

Abstract
How has the credibility revolution reshaped political science? We address this question by using a large language model to classify 91,632 articles published between 2003 and 2023 across 174 political science journals, focusing on causal research designs, transparency practices, and citation patterns. Design-based studies—research strategies that explicitly a research design and the assumptions required for causal identification—have become increasingly common, displacing regression-based analyses that rely primarily on modeling assumptions. Yet as of 2023, studies without an explicit identification strategy still constitute nearly 40% of empirical quantitative work. Within design-based research, survey experiments dominate, while field experiments and quasi-experimental approaches have grown more modestly. Transparency practices such as placebo tests and power analysis remain rare. Design-based studies are concentrated in top journals and among authors at highly ranked institutions, and enjoy a persistent citation premium. The credibility revolution has meaningfully reshaped the discipline, though unevenly and incompletely.
Introduction
Over the last two decades, political science has been swept by what many describe as the credibility revolution: an intellectual movement advocating a change in how researchers conceptualize and conduct causal empirical scholarship (Angrist and Pischke 2010; Dunning 2012; Gerber and Green 2012; Morgan and Winship 2014; Samii 2016; Aronow and Miller 2019; Blair, Coppock and Humphreys 2023; Slough and Tyson 2024; Imbens and Xu 2025). Or has it? In this paper, we examine whether and to what extent the credibility revolution has altered the landscape of empirical causal research in political science.
As an epistemic movement that reshapes what researchers view as credible evidence when assessing empirical, especially causal, claims, the credibility revolution carries genuine transformative potential. Its intellectual core is a reconceptualization of causality. It encourages a shift away from relying primarily on model-based strategies that impose strong functional form assumptions and often implicit views of treatment assignment, toward design-based approaches that exploit how treatments are assigned and rest on explicit identification assumptions. This movement views credibility not solely as a consequence of adopting specific designs such as experiments or natural experiments; additionally, it proposes a clearer articulation of identification assumptions and supporting evidence through transparency practices such as pre-analysis plans, power analyses, and placebo tests (Sekhon 2009). At the same time, critics have raised concerns about external validity and over-reliance on as-if-random designs, underscoring that research design and substantive relevance must be considered jointly. Whether, on balance, this movement has produced an epistemic revolution in political science depends on how widely and deeply these ideas have reshaped actual research practice.
Quantifying the depth and breadth of this potential transformation presents two measurement challenges. The first is scope. What constitutes a sufficiently representative sample of political science to allow statements about the discipline as a whole? Previous research examining methodological changes in other social science disciplines has focused on elite journals or publishing outlets (e.g., Currie, Kleven and Zwiers 2020; Goldsmith-Pinkham 2024). However, the discipline extends far beyond these outlets, which may publish systematically different research than the rest of the field. Consequently, systematic assessment requires analyzing thousands of papers across multiple journal tiers over a sufficient time period to capture potential changes. We address this by analyzing 91,632 articles from 174 political science journals spanning 2003–2023 using large language models (LLMs). This represents, to our knowledge, the most comprehensive analysis of political science research assembled to date.
The second is operationalization. The credibility revolution is an intellectual movement that seeks to transform quantitative empirical research. Translating such ideas into concrete research decisions is left to individual researchers. What changes should we expect to observe if the revolution has taken hold? We focus on three dimensions where such transformation should be evident: the methodological approaches employed in empirical quantitative political science papers, the adoption of practices that enhance credibility beyond research design, and the professional rewards these approaches receive relative to the approaches they aim to substitute.
We leverage advances in LLMs to classify our corpus along these dimensions. The strategy first maps papers by methodology and research goals, then focuses on explanatory empirical quantitative papers, the subset most likely to be shaped by the credibility revolution. Within this subset, we distinguish between design-based approaches that embody the principles of the credibility revolution and traditional model-based approaches, including linear regression and more complex statistical models. We also code the use of ancillary transparency practices, such as pre-analysis plans, power analysis, placebo tests, and robustness checks.
We first assess the extent to which design-based empirical research has substituted model-based approaches in quantitative studies that aim to explain social phenomena, many of which make explicit causal claims. We find substantial—though uneven—substitution. Design-based methods increased from 15% to 40% of explanatory quantitative studies between 2003 and 2023, while model-based approaches declined from 57% to 39%. Yet this rise does not amount to a wholesale replacement. It is driven largely by survey experiments, which accounted for 44% of all design-based papers in 2023, and it remains concentrated in top journals. Overall, the methodological shift is real but circumscribed: traditional strategies—often relying on assumptions that remain unstated or untested—continue to play a central role in explanatory political science research.
Second, we examine the adoption of practices that enhance credibility beyond research design. Such transparency practices remain uncommon even among design-based studies. Pre-analysis plans and power analyses appear in only 6% and 16% of experimental papers, respectively, and placebo tests are reported in just 22% of design-based observational studies. These patterns indicate that the credibility revolution’s influence beyond methodological substitution remains limited and has not yet generated broader shifts in research standards.
Third, we assess whether design-based research receives different professional rewards than model-based approaches. Descriptively, we find evidence of a citation premium for design-based studies: such papers receive more citations than model-based papers, and this gap has widened over time. The premium has grown but now appears to have stabilized, suggesting that the discipline’s reward structure has shifted to favor more credible methods and may be reaching a new equilibrium.
Overall, we find evidence of a partly successful revolution. Design-based methods have grown while model-based approaches have declined, indicating an increase in credible quantitative research. Yet the transformation is neither wide nor deep. It is not wide for two reasons. First, much of political science lies outside the domain where the credibility revolution’s lessons most directly apply: quantitative studies that seek to identify causal relationships. This is not inherently problematic, as high-quality descriptive research remains essential (De Kadt 2025). Second, traditional model-based approaches remain deeply embedded in researchers’ practices. It is not deep in the sense that, even among scholars adopting design-based methods, broader transparency practices remain rare, and the shift is driven largely by one approach–survey experiments, which accounted for nearly half of all design-based papers in 2023. Thus, the credibility revolution in political science is best understood as a reform: a real but limited transformation concentrated in relatively small segments of the discipline.
Our paper builds on Currie, Kleven and Zwiers (2020) and Goldsmith-Pinkham (2024), who quantify the credibility revolution in economics using papers published in elite journals and National Bureau of Economic Research (NBER) working papers. A key limitation of these studies is that NBER affiliates do not represent the economics discipline as a whole. By contrast, we analyze articles from all political science journals with an impact factor above one, allowing us to examine the field well beyond elite institutions. We also extend Garg and Fetzer (2025), who use LLMs to identify causal claims in economics, by systematically classifying design choices and transparency practices across an entire discipline. Finally, our work contributes to a broader literature on how the credibility revolution has shaped political science, including patterns of knowledge production (Grossman, Dinneen and Torreblanca 2025), graduate training and hiring, and the evolution of methodological norms.
The Credibility Revolution and Political Science
We use the term “credibility revolution” to refer to the movement in the social sciences that sought to change how empirical causal research is conducted and evaluated. The motivating ideas marked a clear departure in how causality is conceptualized, empirically investigated, and defended.
The credibility revolution began in applied microeconomics in the early 1990s, driven by natural experiment studies that exploited institutional rules, policy changes, and quasi-random variation to identify causal effects (e.g., Card and Krueger 1994; Angrist and Krueger 1991). These empirical innovations spurred renewed emphasis on, and further development of, the potential outcomes framework and related causal theories.
Before this shift, much quantitative causal research in economics and political science relied on parametric regression models that treated causal and statistical relationships similarly. Researchers typically modeled the conditional expectation of an outcome using treatment variables and covariate (often not clearly separated) and then interpreted regression coefficients as causal effects. Identification assumptions were implicit, often resting on functional-form restrictions or conditional independence claims that were difficult to justify empirically (Angrist and Pischke 2010).
In what follows, we distinguish between two broad approaches to causal inference:
  • Design-based approaches identify causal effects by exploiting a clearly specified research design that provides (or approximates) exogenous variation in treatment assignment and by stating identification assumptions explicitly and separately from functional-form or distributional assumptions.
  • Model-based approaches identify causal effects primarily through parametric or statistical modeling assumptions—such as linearity, additivity, conditional independence, and distributional assumptions—often interpreting regression coefficients as causal effects without an explicit research design or separately articulated identification strategy.
One theoretical underpinning of the credibility revolution is the Neyman–Rubin potential outcomes framework (Rubin 1974; Holland 1986; Imbens and Rubin 2015). In this framework, causal effects are defined as contrasts between a unit’s potential outcomes under different treatment conditions, and causal inference is framed as a missing-data problem. A complementary development is the structural causal model and directed acyclic graph framework, which provides a graphical representation of causal assumptions and identification strategies (Pearl 2009). A central contribution from this literature is the local average treatment effect framework (LATE), which clarifies what instrumental variables identify and under what assumptions (Imbens and Angrist 1994).
Design-based approaches identify causal effects by exploiting randomization, institutional rules, policy cutoffs, or timing that generate exogenous variation and justify explicit identification assumptions (Imbens and Rubin 2015). Natural experiments, randomized controlled trials, instrumental variables, regression discontinuity designs, and difference-in-differences designs are canonical examples. In these designs, assumptions required for identification are stated explicitly, tied to empirical context, and conceptually separated from functional-form or distributional assumptions.
Two designs played especially influential roles. Regression discontinuity, first proposed by Thistlethwaite and Campbell (1960) and formalized for modern use by Lee and Lemieux (2010), became a central tool for credible analysis near institutional cutoffs. Difference-in-differences emerged as another pillar (Card and Krueger 1994). Although the parallel trends assumption can be seen as a functional form restriction, it can also be viewed as an ignorability condition applied to first differences or demeaned outcomes. We treat difference-in-differences as design-based because its use requires researchers to state the research design and identification assumptions clearly.
Political science followed a similar trajectory. Foundational work on field experiments (Gerber and Green 2000, 2012) and the synthesis of natural experiments in Dunning (2012) helped institutionalize design-based reasoning and advance credibility norms across the discipline. Because of the central role of electoral studies, regression discontinuity became a powerful and widely used tool (Eggers et al. 2015). LATE remains central to modern design-based instrumental-variable analysis in political science (Sovey and Green 2011; Lal et al. 2024). Difference-in-differences is also widely applied, though recent work highlights challenges posed by heterogeneous treatment effects, staggered adoption, and treatment reversals, all common in political science settings (Chiu et al. 2025). Together, these developments reshaped what counts as credible causal evidence in the field.
The emphasis on credibility extends beyond research design. Robustness and sensitivity analyses—placebo tests, balance checks, manipulation tests, and formal sensitivity procedures—assess how results respond to departures from identifying assumptions (Sekhon 2009; Blair, Coppock and Humphreys 2023; Cinelli and Hazlett 2020). At the same time, transparency practices such as pre-analysis plans, power analyses, open data and code, and clear reporting of deviations from pre-registered plans have become increasingly common. Journals and professional associations have institutionalized these norms through data and code policies and reproducibility checks, embedding transparency as a central component of credible empirical practice.
Debates accompany this shift. Critics argue that design-based estimands can be unduly local or may emphasize identification over substantive relevance (e.g., Heckman and Smith 1995) or external validity (Findley, Kikuta and Denly 2021). Others question strong assumptions behind randomized-control trials or as-if random designs (e.g., Deaton and Cartwright 2018). These debates shape how credibility is interpreted across fields and underscore that research design and theoretical relevance are complementary.
Quantifying the impact of the credibility revolution.
The core ideas of the credibility revolution generate testable empirical predictions about how quantitative explanatory research should change. We focus on three dimensions: (a) methodological practices, (b) transparency and research standards, and (c) professional incentives.
First, if the credibility revolution has taken hold, design-based methods should increasingly displace purely model-based approaches in quantitative explanatory research, with growth spread across multiple design-based strategies.
Second, the credibility revolution should lead to greater use of transparency practices that strengthen the credibility of causal claims. Pre-analysis plans, power analyses, placebo tests, and transparent reporting of data and code should become more prevalent, as each addresses a distinct but complementary credibility concern.
Third, the credibility revolution implies a shift in the standards by which causal research is evaluated, which in turn should be reflected in disciplinary rewards. If the field increasingly views design-based approaches and transparency practices as markers of credible inference, then research employing these strategies should attract greater attention. As a result, design-based studies would be expected, on average, to receive more citations and to be published more frequently in prominent outlets than otherwise comparable model-based work.
Data and Measurements
We explore these predictions using a corpus of political science research from the past two decades. In this section, we describe the data, key measurements, and methods used to construct these measures.
Data.
We begin with a set of 188 journals identified by Clarivate as part of the political science discipline and having a Scientific Journal Rankings (SJR) score of at least one. This threshold roughly corresponds to the upper tier of the field in citation influence and focuses attention on outlets that sit near the center of political science’s citation network. From this list, we removed journals for three distinct reasons: indexing (three journals not indexed in Scopus), quality control (three journals that lacked peer review), and language or format (seven journals that were not published in English and one book series). The final sample contains 174 English-language, peer-reviewed, Scopus-indexed political science journals. Our aim is not to cover every outlet, but to assemble a comprehensive corpus of articles that shape debates in the discipline.
Figure 1: Number of papers published each year in our full sample from 2003 to 2023 ($n=91,632$) and by subfield, with loess fits. The red line shows total papers; blue, American Politics; gray, Comparative Politics; and green, International Relations.
Using Scopus, we extracted journal-level metadata, including citation statistics, and article-level information for all publications in the selected journals. This process produced a dataset of 129,751 articles published between 2003 and 2023. To systematically code research design choices, we obtained full-text versions of the articles where possible and successfully retrieved the complete text for 91,632 articles from 156 journals. Full-text availability varies somewhat across publishers and subfields, so the full-text corpus is broad but not exhaustive. We then used a combination of supervised machine learning and a commercial LLM, gpt-4o, to classify articles along 19 research design dimensions (see description of the approach below). Figure 1 shows the yearly number of published papers in our full-text sample, both overall and by subfield. The volume of published work has grown rapidly. From 1,546 papers in 2003, the count has more than quadrupled to over 8,109 in 2023.
Key variables.
We construct a series of variables to characterize each paper’s methodological features, substantive focus, and transparency practices. Details of the coding procedure appear in the Supplementary Materials.
We first assign each paper a subfield label drawn from six categories: American Politics, Comparative Politics, International Relations, Methodology and Formal Theory, Political Theory and Philosophy, and Public Policy/Administration. The label reflects the paper’s primary substantive contribution.
We then identify whether the paper is an empirical quantitative study. A paper is coded as empirical quantitative if it conducts its own analysis of observational or experimental data, including reanalyses of existing datasets. Simulation exercises or purely methodological demonstrations without data are not included. Papers whose main contribution is qualitative, formal-theoretic, or normative are coded as non-quantitative for our purposes, though we distinguish these categories in the underlying coding. We assign general goals only to empirical quantitative papers, because only these studies contain measured relationships between variables that can be classified consistently as descriptive, explanatory, or predictive.
Explanatory papers investigate the causes or consequences of social phenomena and develop evidence consistent with one or more causal relationships, even if they do not always use explicit causal language. Descriptive papers center their contribution on measurement or characterization, for example by introducing new indicators, establishing baseline levels, or documenting group differences, without developing or testing a causal mechanism. Predictive papers focus on forecasting an outcome or identifying variables that best predict it, typically using out-of-sample or cross-validated performance as an evaluation criterion, and do not interpret those variables as causes. In cases where papers combine elements of these goals, we code the general goal based on the primary contribution. Because explicit causal designs are primarily used in explanatory work, we restrict our coding of primary research design to papers classified as explanatory.
Figure 2: Sequential classification of type of studies. Flows trace papers from the full sample to quantitative studies, then to descriptive, predictive, and explanatory goals, and finally to design-based versus model-based and other approaches. The percentage numbers use the full sample of studies, $91,632$, as the denominator.
Figure 2 illustrates this sequential classification. It reveals three patterns from our classification. First, quantitative, empirical papers are a minority. They represent 41% of all papers pooled across years. Second, within quantitative work, explanatory papers dominate (81%). Descriptive papers constitute only 18% of empirical quantitative papers. Predictive papers remain rare (240 across the entire period). Third, among explanatory studies, model-based methods remain the majority. Design-based methods account for 29% of explanatory papers, suggesting substantial but still limited methodological transformation.
Averages over two decades may mask important temporal dynamics. Figure 3 shows that the share of empirical quantitative papers has increased steadily, reaching about 48% by 2023. Disaggregation by subfield reveals similar upward trends but persistent differences in baseline levels: In 2023, 27% of International Relations papers use quantitative methods, compared to 57% in Comparative Politics and 84% in American Politics.
Figure 3: The proportion of empirical quantitative papers overall and by subfield. The red curve shows the overall share, and the blue, gray, and green curves show trends for American Politics, Comparative Politics, and International Relations. Lines are lowess fits with 95% confidence intervals; points indicate raw annual proportions.
For the $30,296$ explanatory empirical quantitative studies, which are the main focus of our analysis in the next section, we classify each paper’s primary research design into more than a dozen categories spanning design-based approaches model-based approaches, and other methods. Design-based approaches include experimental studies—field, survey, and lab experiments—and a range of non-experimental identification strategies. These non-experimental methods include (1) methods based on unconfoundedness and overlap, such as matching and reweighting techniques; (2) natural and quasi-experiments; [^1] These studies use exogenous variation in treatment not controlled by the researcher, such as shocks or sudden institutional rules (e.g., natural disasters, draft lotteries, externally implemented policy changes). They fit within unconfoundedness designs, but researchers often analyze them with regression methods without explicitly stating unconfoundedness or selection-on-observables assumptions. (3) instrumental variables designs; (4) regression discontinuity and regression kink designs; and (5) difference-in-differences designs; [^2] This category includes broadly defined causal panel analysis under parallel trends (Chiu et al. 2023). Its implementations include canonical difference-in-differences, two-way fixed effects models, event studies, longitudinal approaches to staggered treatment adoption, and related variants. All assume some version of the parallel trends assumption. (6) the synthetic control method; and (7) other observational strategies that attempt causal inference by explicitly specifying an assignment mechanism and identifying assumptions such as unconfoundedness, parallel trends, continuity, or exclusion restrictions.
Model-based approaches estimate associations using parametric models without an explicit design component, such as linear regression, generalized linear models, structural equation models, and time-series models including ARMA, GARCH, and vector autoregression. Other methods form a residual category, capturing approaches not covered by the pre-specified design-based or model-based categories. Examples include network analysis, text-as-data methods, agent-based modeling, and purely correlational analyses such as simple bivariate correlations and crosstabs.
Although linear regression can recover causal effects in both experimental and non-experimental settings (Angrist and Pischke 2009; Aronow and Samii 2016; Imbens and Xu 2025), we classify a study as design-based only when it explicitly identifies the source of causal variation, such as randomized treatments or shock-induced variation in natural or quasi-experiments. Many regression-based studies in the corpus do not do this. Instead, they follow what is often called a “kitchen sink regression” approach, either failing to separate the treatment from covariates or adding many covariates in the hope of avoiding omitted-variable bias, without articulating an identification strategy distinct from the model specification. We classify these studies as model-based. In fact, a vast majority of model-based studies in the corpus, 71%, rely on simple regression analysis of this form.
For these explanatory studies, we also code whether the paper states key identification assumptions, defined as explicit statements of the core assumption underlying the design (for example, unconfoundedness, parallel trends, continuity, and exclusion restrictions); merely naming a method is not sufficient. We then record whether the paper makes a causal claim. Explicit claims use terms such as “causes,” “effect,” or “impact.” Implicit claims are coded when the paper frames its contribution in terms of causes and consequences or interprets coefficients as causal, even if it avoids direct causal terminology.
Finally, we code several features of the paper and its transparency practices, including the sample size in the main analysis, whether authors fielded original surveys, whether the paper reports placebo tests, and whether it includes a power analysis. We focus on these practices because they appear across designs and are widely used to assess credibility. These variables allow us to document methodological choices, causal reasoning, and transparency practices across the quantitative political science literature.
An important caveat is that we measure these variable by taking at face value what others have reported in their papers, not by assessing how well they executed a particular research design or whether their reports contain mistakes or misrepresentations of methods. Recent replication studies of political science research have found that, across various designs, there remain many problems with the validity of assumptions, statistical power, and errors in execution (e.g., Stommes, Aronow and Sävje 2023; Lal et al. 2024; Chiu et al. 2025).
LLM pipeline and human validation.
Having defined our measures, we next describe how we coded them at scale. We developed an LLM-based coding framework and conducted a structured human validation to generate reliable measures from our corpus of political science articles. We first gathered the raw text of all scraped papers and truncated any article exceeding 50,000 tokens (about 35,000 words) to manage cost and ensure processing efficiency. In practice, this affected only a small number of long review articles, and truncation typically removed appendices rather than main empirical content. We then sent each paper’s text, together with a custom prompt, to OpenAI’s gpt-4o model via the batch API. The System Prompt outlined 19 research dimensions to code, including research design, transparency practices, and causal claims (see Supplementary Materials). Using OpenAI’s structured outputs, we constructed a JSON schema to ensure consistent formatting, and the model coded each article according to this schema. To refine the prompt, we iteratively tested eight versions by qualitatively evaluating outputs for a sample of 116 papers. Each author read a couple dozen papers, noted errors, and proposed revisions. After eight rounds, we finalized the prompt and evaluated it on the five key variables.
We validated five core variables with human coders that anchor our main analysis: (1) whether a paper is an empirical quantitative study, (2) its subfield, (3) its primary research design, (4) whether it states key identification assumptions, and (5) whether it makes a causal claim. Subfield and research design are multi-label variables spanning six and seventeen categories; the remaining variables are binary. Task 1 evaluated whether papers were empirical and quantitative using a random sample of 200 papers. Task 2 assessed the remaining variables using another 200 papers identified by the model as empirical and quantitative, with one paper removed after review. Four research assistants coded each paper in pairs, with disagreements resolved by a third coder in consultation with an author. Although modest relative to the corpus, these samples are sufficient to estimate accuracy with useful precision, and additional spot checks across journals yielded similar patterns. Accuracy was high: Task 1 reached 98% agreement with human coders. In Task 2, accuracy was 83% for subfield, 73% for research design (84% among design-based studies), 78% for identification assumptions (90% among design-based studies), and 82% for causal claims (87% among design-based studies).
Overall, these exercises show that LLMs can generate reliable large-scale measures of methodological features in political science when supported by targeted human checks. The LLM performs notably better among design-based studies because these papers tend to have more structured methods sections, which affects coding difficulty.
Main Findings
To evaluate how far and how deep the credibility revolution has reshaped quantitative empirical research in political science, we present descriptive evidence on the scope and pace of methodological change across the discipline. We examine shifts in research designs, empirical practices, and differences in scholarly influence between studies adopting design-based methods and those using traditional approaches.
Design-Based Methods Rose Steadily But Unevenly
How have methodological practices shifted in quantitative explanatory research? If the credibility revolution has taken hold, design-based strategies should increasingly displace model-based approaches. Figure 4(a) traces these aggregate trends from 2003 to 2023. Model-based methods (gray) dominated for most of the period, but design-based approaches (black) rose from 15% to 40%. By 2023, the two approaches accounted for nearly equal shares of explanatory quantitative research, while other methods, a residual category that includes analyses without explicit designs, constitute 18%.
(a) Design-based vs. model-based studies
(b) Design-based studies by subfield
Figure 4: Trends in the use of design-based methods among explanatory empirical quantitative papers.
Within this overall shift, subfield differences are relatively modest. As Figure 4(b) shows, American Politics began slightly ahead in 2003 and widened its lead through the mid-2010s, but by 2023 all three major subfields show substantial adoption: American Politics at 52%, Comparative Politics at 38%, and International Relations at 43%.
To understand which specific designs drive this growth, Figure 5 breaks down the composition of design-based work. Survey experiments account for much of the increase and become the single most common design by 2023. Difference-in-differences and matching or reweighting methods based on unconfoundedness also expand, while instrumental variables, field experiments, and regression discontinuity designs maintain smaller but steady shares. Natural/quasi-experiments and synthetic control remain rare.
Next, we examine how each method has evolved over time. Figure 6 traces these trajectories for both experimental and non-experimental approaches. Figure 6(a) reports trends for the three experimental approaches. Survey experiments grow rapidly, rising from about 4% of all explanatory quantitative studies in 2003 to more than 15% by 2023, and constitute roughly 45% of all design-based research by the end of the period. Field experiments remain uncommon but show a modest upward trend. By contrast, lab experiments decline sharply after 2016. This decline likely reflects the increasing availability of online and survey-based experimental platforms, which allow researchers to conduct experiments at lower cost, with larger and more diverse samples, and without the logistical constraints of in-person laboratory settings.
Figure 5: Relative share of design-based studies in all explanatory empirical quantitative studies, by year.
Figure 6(b) reports trends for six non-experimental design-based methods. Difference-in-differences (including its various implementations) grows almost monotonically and reaches more than 6% of all explanatory empirical quantitative papers in 2023. Matching and reweighting methods increase to about 5% around 2012 before plateauing through the end of the period. Instrumental variable designs rise from 3% in 2003 to nearly 5% in 2012, then gradually decline over the past decade. Regression discontinuity designs increase slowly but remain near 2% in 2023, reflecting limited opportunities for such designs. Natural and quasi-experiments remain around 1% of explanatory empirical quantitative studies throughout. Synthetic control appears around 2012 and increases modestly in recent years, but still constitutes less than 1% of explanatory empirical quantitative papers.
(a) Experimental
(b) Non-experimental
Figure 6: Popularity of design-based methods by year. The denominator is the total number of explanatory empirical quantitative studies in each year.
Taken together, these patterns show that the credibility revolution has altered methodological practice, but unevenly across designs. Survey experiments account for most of the overall shift, followed by difference-in-differences designs, while other design-based approaches have grown only modestly or remained niche.
Design-based studies are concentrated in influential journals.
We found that design-based methods have risen slowly but steadily over the last two decades, reaching parity with traditional model-based approaches only in 2023. This pattern suggests that credibility-revolution practices have diffused at a modest pace and remain far from universal. One concern, however, is that treating all articles equally may mask differences across publishing venues. More influential journals may adopt emerging methodological standards earlier and thereby shape broader disciplinary expectations. If so, equal weighting could understate both the extent and the speed of change occurring in more widely read higher-impact outlets.
To assess this possibility, Figure 7 plots the share of design-based and model-based studies among explanatory quantitative papers when weighting each article by the SJR score of its publishing journal. The left panel presents trends for the 20 highest impact factor journals; the right panel presents all others. In higher-impact journals, design-based approaches rise more quickly and surpass model-based work several years earlier than in the rest of the field. In contrast, model-based methods remain predominant in lower-impact outlets, although design-based work also increases over time. These patterns indicate that the credibility revolution has diffused unevenly across publishing venues, with faster uptake in journals that play a central role in shaping methodological norms in the discipline.
(a) Studies published in top 20 journals
(b) Studies published in other journals
Figure 7: Proportion of design-based and model-based studies by journal ranking. The denominator is the total number of explanatory empirical quantitative studies published each year, weighted by SJR (a measure of journal impact). Observations are weighted by SJR. The left panel shows papers in the top 20 journals; the right panel shows papers in other journals. Curves with 95% confidence intervals are lowess fits.
(a) Experimental
(b) Non-experimental
Figure 8: Proportion of each design-based method by journal ranking and year. The denominator is the total number of explanatory empirical quantitative studies published each year. Top 20 journals are shown in blue; all other journals in red.
Impact-factor weighting adjusts for differential influence across papers but does not reveal where those differences originate. Figure 8 shows which specific design-based methods are driving the methodological shift and how these patterns differ between top journals and the rest of the field. It compares the prevalence of each method in the top 20 journals versus all other outlets. With most design-based methods, top journals adopt design-based strategies earlier and at higher rates, indicating that influential outlets play a disproportionate role in shaping the trajectory of causal inference practices in the discipline.
Two broad patterns emerge. First, survey experiments are the principal driver of growth in design-based work. They rise sharply in top journals—reaching more than 25% of explanatory quantitative papers by 2023—and show steady, if slower, growth in other outlets. Difference-in-differences designs also expands consistently, with faster growth in top journals. Field experiments and regression discontinuity designs are substantially more common in top outlets, even though their overall frequencies remain modest. Lab experiments decline in both tiers but remain slightly more prevalent in top journals.
Figure 9: Proportion of design-based studies among top 30 political science journals (by SJR score) with 95% confidence intervals. The numerator is the number of design-based studies published between 2003 and 2023. The denominator is the total number of explanatory empirical quantitative papers published during the same period. Journals ranked as top-10 by SJR are painted in black.
Second, several methods exhibit limited or uneven diffusion. Matching and reweighting methods peak around 2015 and then slightly declined in top journals. Instrumental variable designs rise and fall in a similar pattern across journal tiers, though always at higher levels in top outlets, and peaked in more recent years. Natural and quasi-experiments remain rare everywhere, while synthetic control shows only recent and modest growth. Overall, the figure highlights that the credibility revolution has been propelled largely by a small set of design-based methods, especially survey experiments and difference-in-differences, with high-impact journals consistently at the leading edge of this shift.
Figure 9 shows the proportion of design-based studies among the top 30 political science journals between 2019 and 2023. Journals in the top 10 generally publish a higher share of design-based work, reflecting their stronger embrace of credibility-revolution practices. One notable exception is Political Analysis, which appears lower despite being a top-10 journal by impact factor. This is expected: as a methods journal, it publishes many papers developing statistical or computational models rather than applying design-based approaches to empirical problems. Overall, the figure illustrates that the most prestigious outlets tend to lead the shift toward design-based methods, while journals further down the ranking show more modest adoption.
Authors from higher-ranked institutions are more likely to adopt design-based methods.
Results thus far show that design-based methods were adopted earlier—and are more common–in high-impact influential journals than in other outlets. This pattern raises the possibility of uneven diffusion across the discipline. To explore this, we examine authors’ institutional affiliations as an additional indicator. Institutional rank offers a complementary perspective on whether uptake has varied across different segments of the discipline.
Figure 10: Adoption of design-based methods by authors’ average institutional rank. Lower ranks indicate more prestigious institutions. The sample includes all papers with institutional affiliation data matched to Shanghai rankings. Circles represent means of 20 bins, and the lines show lowess curves with 95% confidence intervals.
We use the Shanghai Academic Ranking of World Universities (ARWU) to rank institutions. We extracted institutional affiliations for all authors and matched 77,123 author–institution pairs to Shanghai-ranked universities across 59,248 papers, covering 65% of our full-text sample. For each paper, we compute the average rank of the authors’ institutions. Details on the matching procedure appear in the Supplementary Materials.
Figure 10 shows design-based method adoption by institutional rank. Among authors from the top 40 institutions, adoption rates decline gradually—from roughly 50% at rank 1 to about 28% at rank 40—and then level off around 10% for lower-ranked institutions. These patterns suggest that the methodological shift associated with the credibility revolution has been more pronounced among authors at highly ranked institutions.
This concentration parallels the journal-tier patterns. Whether it reflects early uptake that will diffuse more widely or a more persistent stratification remains uncertain. What is clear is that even among highly ranked institutions, the shift has been gradual: design-based methods took nearly two decades to reach parity with model-based approaches. Moreover, within design-based work, the rise has been driven primarily by survey experiments rather than broad adoption across multiple design-based strategies.
Credibility-Enhancing Practices Progressed Modestly
Research design is not the only source of credibility in quantitative causal inference. Scholars also rely on empirical practices that strengthen the plausibility of their claims. Pre-analysis plans, for instance, constrain researcher discretion and reduce the risk of $p$-hacking. Original survey data collection enhances credibility by enabling precise measurement of key variables and confounders, minimizing measurement error and omitted variable bias. Placebo tests further demonstrate that identifying variation operates through the theorized mechanism rather than spurious channels. Power analyses similarly bolster credibility by assessing whether a study is adequately equipped to detect meaningful effects, thereby reducing the likelihood of Type II errors and clarifying the evidentiary value of null findings. Importantly, these credibility-enhancing tools often complement but are not limited to design-based approaches: researchers using model-based methods can, for example, employ placebo tests (e.g., showing that treatment does not affect prior outcomes) or use simulations to assess statistical power.
We examine six dimensions of empirical practice used to enhance the credibility of causal claims: (1) sample size, (3) original survey data collection; (3) placebo tests, (4) explicit statement of identification assumptions, (5) pre-analysis plans, and (6) power analysis. [^3] We also explored other criteria, such as clearly stating the causal quantity of interest, but we found it is highly correlated with the adoption of design-based methods. In other words, it is very rare to see researchers using model-based methods to define such quantities explicitly. We assess the prevalence of these practices separately for experiments, design-based observational studies, and model-based approaches. This disaggregation reflects the fact that disciplinary norms impose some credibility practices more strongly on experiments than on observational research, and pooling them would obscure meaningful differences in trends.
Figure 11 reports results for sample sizes, original data collection, placebo tests, and explicit statements of identification assumptions. Sample sizes increased substantially across all three categories. Experiments had median sample sizes of 490 in 2003, growing to 1,601 by 2023. The number of observations in studies with model-based methods grew from 984 in 2003 to 2,059 by 2023. For design-based observational studies, the number started at 784 in 2003 and reached 3,019 by 2023, showing the steepest growth. All else equal, these increases imply better-powered research regardless of method.
Figure 11: Transparency and credibility practices over time for quantitative explanatory papers. Top left: median sample size. Top right: proportion using original survey data. Bottom left: proportion conducting placebo tests. Bottom right: proportion explicitly stating identification assumptions. Experimental, design-based nonexperimental, and model-based studies are in blue, red, and gray, respectively.
As expected, experimental research relies heavily on original data collection, with nearly 75% of experimental papers reporting original data collection by 2023. In contrast, design-based observational research and model-based work remain flat below 9% throughout the study period, suggesting that these papers typically rely on existing datasets.
Placebo tests remain concentrated in design-based observational research. This makes sense given these methods rely on design features of the data-generating process for identification rather than generating identifying variation directly as in experiments (Eggers, Tuñón and Dafoe 2024). Their use grows from under 5% in 2003 to roughly 20% by 2023, but levels remain modest even within the category where they are most relevant. Placebo tests remain rare in both model-based and experimental work, despite their usefulness for probing alternative explanations or assessing identifying assumptions.
Explicit statements of identification assumptions show a more encouraging pattern. They have become nearly universal in design-based work, both observational and experimental, surpassing 90% by 2023. Model-based papers also reach over 60%, though without strong upward trends. The discipline has become substantially more transparent about the assumptions required to interpret results causally, although the assumptions associated with model-based methods are usually more difficult to interpret and verify.
Figure 12: Best practices for quantitative explanatory papers. Left: proportion reporting pre-analysis plans. Right: proportion reporting power analysis. Experimental, design-based nonexperimental, and model-based studies are in blue, red, and gray, respectively.
Pre-analysis plans show sharp growth but almost exclusively within experimental work, reaching approximately 35% by 2023. Design-based observational studies exhibit minimal adoption, rising from essentially zero to roughly 3%. Power analyses follow a similar pattern: growing from 4% in 2003 to only 15% in 2023. Power analysis can inform effect size interpretation and null result credibility across research designs, yet it remains rare even in experimental settings where it most commonly appears.
These patterns indicate the credibility revolution has produced an uneven transformation in research practices. Explicit identification assumptions have diffused widely and are now standard practice in design-based research. Other transparency practices, however, remain concentrated in specific method or show limited adoption overall. The credibility revolution has altered how researchers articulate identification strategies, but has not reshaped research standards comprehensively.
Citation Premium for Design-Based Methods Persists
A discipline-wide transformation requires not only shifts in research design choices, but also changes in what the discipline values and rewards. Citations, despite their limitations, offer one useful proxy: they reflect what gets read, built upon, and treated as influential. Because citation counts affect promotion, hiring, and professional recognition, they provide insight into how the field allocates prestige.
If the credibility revolution reshaped political science, design-based studies should receive more citations than model-based work. We refer to this difference—the average number of additional citations that design-based papers receive compared to model-based papers published in the same year—as the citation premium for design-based studies. A key limitation of this analysis is that we do not account for topic, subfield, journal placement, or research quality, so these patterns should not be interpreted as causal effects.
(a) Citation count in 2023 by publication year
(b) Citation premium by publication year
Figure 13: Citation premium for design-based methods. The left panel shows the average 2023 citation count for design-based (blue) and model-based (red) studies published in each year. The right panel shows the citation premium (the difference in average citations between design-based and model-based studies) measured relative to the publication year for papers published in 2003, 2007, 2011, 2015, and 2019.
Comparing citation levels across years is complicated by the fact that citations accumulate over time, so more recent papers have had fewer years to be cited, and design-based methods have become more common in later cohorts. To address these issues, we examine both raw citation counts by publication year and cohort-specific citation trajectories.
Figure 13(a) presents the average number of citations received by design-based and model-based papers. The citation premium emerges in the early 2000s, grows through the late 2000s, peaks around 2011, and appears to narrow in more recent years. Two considerations likely contribute to this pattern. First, newer cohorts have had less time to accumulate citations, mechanically reducing observable gaps. Second, as design-based methods became more common—including in influential journals—the bar for publishing model-based work may have risen, producing stronger model-based papers in recent years and narrowing the citation differential. In addition, the broader diffusion of design-based approaches means that some studies may adopt these methods uncritically without fully meeting their underlying assumptions, limiting their potential to generate a citation advantage.
Figure 13(b) shows the citation premium for cohorts of papers published in 2003, 2007, 2011, 2015, and 2019, measured relative to each paper’s publication year. The premium is consistently positive, but varies by cohort. The 2003 cohort shows a small advantage of about seven additional citations after twenty years. The 2007 cohort rises more quickly, reaching roughly fifteen extra citations after sixteen years, a pattern matched by the 2011 cohort within eleven years. The 2015 cohort accelerates further, gaining about ten additional citations within eight years, compared to seven for the 2011 cohort at the same point. The 2019 cohort tracks the 2015 pattern closely, suggesting that the premium may have stabilized. Overall, design-based papers continue to receive more citations than comparable model-based work, but the growth of this premium appears steady rather than increasing over time.
With important caveats in mind, including potential differences in average research quality, our results suggest that the credibility revolution has been rewarded in the discipline, though unevenly across cohorts. At the same time, the recent stabilization indicates that the advantage of design-based work may now be leveling off.
Discussion
Our analysis paints a mixed picture of the credibility revolution in political science. On the one hand, design-based methods have moved from the margins to rough parity with traditional model-based approaches in quantitative explanatory research. Design-based strategies now account for about 40% of such papers, and they receive a persistent citation premium relative to model-based work. On the other hand, this transformation is not universal and remains highly stratified. It is disproportionately driven by a single design—survey experiments—and is concentrated in top journals and among authors affiliated with highly ranked institutions. Transparency-enhancing practices beyond research design, such as pre-analysis plans, power analyses, and placebo tests, remain far from universal even among design-based studies. In this sense, the credibility revolution is better described as a substantial but partial reform than as a wholesale reordering of empirical practice.
Several developments are nonetheless welcome. First, the discipline has become less reliant on parametric modeling as the default route to causal claims. The growth of experiments, regression discontinuity, and other designs has shifted attention from functional form assumptions to explicit statements about treatment assignment and identifying variation. Second, researchers are now much more explicit about the assumptions required for causal interpretation: identification assumptions are stated in more than 90% of design-based papers and in a majority of model-based papers. Third, median sample sizes have increased across methodological families, which, all else equal, should improve statistical power and reduce the risk that published findings are driven by small-sample noise. Together, these changes move the discipline closer to the standards articulated by proponents of the credibility revolution.
Simultaneously, the evidence suggests that the credibility revolution is not deep in at least two respects. First, the adoption of design-based methods is highly concentrated in survey experiments, which account for nearly half of all design-based papers toward the end of the period. Other designs central to the credibility revolution playbook—difference-in-differences, regression discontinuity, natural and quasi-experiments, synthetic control—remain rare and often exhibit non-monotonic usage patterns, suggestive of methodological fashions rather than durable practice. This concentration limits the types of questions that can be credibly addressed and leaves many important settings under exploited. Second, outside of identification assumptions, other credibility-enhancing procedures remain weakly institutionalized. Placebo tests are used in only about one-fifth of design-based observational studies and are very rare elsewhere. Pre-analysis plans and power analyses are mostly confined to experimental work and remain the exception even there. If the goal is to make published evidence more diagnostic of causal claims, these ancillary practices must become more routine across methods, not just in a subset of high-profile experiments. Moreover, clear and precise statements of the exact causal quantity of interest remain uncommon even among design-based studies.
The reach of the credibility revolution is also not wide. Design-based methods were adopted earlier and at higher rates in the discipline’s most prestigious journals and among authors at top-ranked institutions. Regression discontinuity designs and field experiments, for example, are disproportionately concentrated in top-10 outlets and in the upper tail of the institutional ranking, whereas adoption rates flatten beyond roughly the top-50 institutions. These patterns are descriptive and likely reflect selection on training, resources, and submission strategies as well as differential “internalization” of credibility principles. They nonetheless raise the possibility of a two-tiered discipline in which standards for credible causal inference vary by institutional location. One implication is that graduate education and methodological training will be crucial for broadening the reach of design-based methods and transparency-enhancing principles. The existence of a citation premium for design-based work and clear expectations in top journals creates aligned career incentives; whether those incentives diffuse to departments with fewer resources will depend on how quickly curricula, summer schools, and methodological communities update outside the current core.
A third implication concerns the relationship between design-based methods and structural or model-based approaches. Structural estimation—estimating structural parameters of a formal model—is extremely rare in political science, and we do not code hybrid designs that combine structural models with design-based identification for this reason. Our reading of the field is that such combinations remain uncommon, in stark contrast to economics, where design-based estimates are increasingly integrated with structural models to study mechanisms, counterfactual policies, and welfare-relevant quantities. This gap matters because many substantive questions in politics—about institutional reform, long-run dynamics, or strategic interaction—require moving beyond average treatment effects in a single setting. Future work could extend our coding to identify papers that embed design-based variation within structural or formal frameworks, and political science more broadly could benefit from a more systematic integration of credible research designs with structural modeling and formal theory.
A fourth implication is that the credibility revolution is reshaping, but not displacing, other modes of inquiry. Quantitative explanatory work is a minority of political science, accounting for about 41% of papers in our corpus, and this share has grown but not overwhelmed qualitative, descriptive, or theoretical work. Our findings are consistent with an emerging equilibrium in which design-based quantitative work coexists with, and is complemented by, qualitative research, archival work, fieldwork, and both formal and normative theory. As standards for causal identification rise in empirical work, the comparative advantages of other approaches may become clearer rather than weaker. Qualitative and theoretical work can help generate and refine hypotheses, elucidate mechanisms, and interpret heterogeneous effects, while descriptive research can map phenomena and measure constructs that design-based studies then exploit. In this view, the credibility revolution is most productive when it promotes a division of labor and mutual reinforcement across methods rather than a zero-sum competition.
Taken together, these results suggest several lessons for the discipline. First, there is room—and need—for more creative application of the full portfolio of design-based methods rather than continued reliance on a few dominant templates. Second, greater integration of design-based identification with structural and formal models would help political science use credible evidence to address counterfactual and mechanism-oriented questions. Third, expanding the use of credibility-enhancing practices such as pre-analysis plans, power analyses, and systematic placebo and sensitivity checks would align everyday research practice more closely with the aspirations of proponents of credible social science. Finally, the evolution of the discipline will depend on building stronger linkages across methodological communities: between design-based empiric and qualitative fieldwork, archival research, and theory-building. Our results show real progress, but they also indicate that the credibility revolution in political science remains an ongoing reform project whose full potential has yet to be realized.